# Bisecting masses with families of parallel hyperplanes

### v1

@@ Abstract.
The general result follows from the case of measures that are supported on a finite set with an odd number of points. The proof for this case is inspired on ideas of differential and algebraic topology, but it is a completely elementary parity argument.
@@
Soberón’s research is supported by NSF CAREER grant DMS 2054419 and a PSC-CUNY TRADB52 award.
@@ 1. Introduction
Understanding how one can split families of measures on real spaces into pieces of the same size is a central topic in topological combinatorics [RoldanPensado2022, Zivaljevic2017]. The classic example, known as the ham sandwich theorem, states that given $d$ smooth probability measures on $\mathds{R}^{d}$, there exists a hyperplane simulteaneously splitting each measure in half.
@@
A myriad of related results have appeared since then, many of them recover the foodie references (cakes [cccg98-akiyama-perfect], spicy chickens [Karasev:2014gi], pizzas [barba2019sharing], etc.), most of these results are proven with topological methods, which range from fixed point and Borsuk–Ulam type theorems to more advanced techniques in algebraic topology (cohomological index theory, equivariant obstruction theory, spectral sequences, etc). Even splitting measures with families of few hyperplanes can lead to very rich topological statements (see, for instance, the history of the Hadwiger–Grünbaum–Ramos theorem [Blagojevic:2018jc]).
@@ 1.1. Our result
A central instance of this definition is when a finite measure $\mu$ is absolutely continuous with respect to the Lebesgue measure (or more generally, $\mu(h)=0$ for every affine hyperplane $h$), in this case the biecting condition is simply:
@@
  • The case $d=1$, $n=1$ and any $k$ is the “necklace splitting theorem”. This was orignal proved by Hobby and Rice [Hobby:1965bh] and then rediscovered in the 80’s with two new proofs [Goldberg:1985jr, Alon:1985cy]. As $S(k,k)=1$, we don’t impose any new requirement on $k$.
@@ 1.4. Summary
The proof of Theorem 1.1 has three steps. First, we construct a particular family of measures supported in sets of an odd number of points for which the number of bisecting arrangements is exactly $N$. Second, we show that for any deformation of these point sets the parity of the number of bisecting hyperplane arrangements is invariant. Finally, we show that the result for measures supported on an odd set of points implies the general result.
@@ 2. The ham sandwich theorem revisited
In this section we show how our method give a new and elementary proof of the ham sandwich theorem. This is a particular case of the general proof described in the next sections, but may help the reader build intuition about this approach.
@@ 3.1. Points in generic position
The proof is analoguous to the one for Lemma 3.1. Let $X=X_{1}\cup\dots\cup X_{k}$. Pick a point $p_{i}\in X_{i}$ for each $i=1,\dots,k$, and consider the $l-2$ vectors of the form $p-p_{i}$ for $p\in X_{i}\setminus\{p_{i}\}$. These, together with $B_{L}$, form $d-2$ linearly independent vectors, which gives us the basis for $K$. The $k$ translates of $K$ containing each of $p_{1},\dots,p_{k}$ are the affine spaces we were looking for. ∎
@@ 3.3. Parity under deformation
then $h^{r}(s_{r}+\epsilon)$ bisects the unbalanced color at time $s_{r}$.
@@ 4. Necessity of conditions
In this case, we construct $(M+1)/2$ pairs of parallel segments as before, and then remove one segment of the final pair. Assume that we have a valid arrangment of hyperplanes that bisects this last segment. If at least two hyperplanes cut this last segment, then we do not have enough hyperplanes to bisect the remaining $(M-1)/2$ pairs of tiny parallel segments. If exactly one hyperplane cuts the last segment, we may assume without loss of generality that is is not one of the hyperplanes orthogonal to $L_{1}$ (otherwise, replace the roles of $L_{1}$ and $L_{2}$ in the rest of the argument). Then, since we still have to use an odd number of hyperplanes orthogonal to $L_{1}$, the argument follows as in case $1$. ∎
$\{(x_{1},x_{2},x_{3}):x_{3}=0,x_{1}=x_{2},x_{1}\geq 0\}$.

### v2

@@ Abstract.
The general result follows from the case of measures that are supported on a finite set with an odd number of points. The proof for this case is inspired by ideas of differential and algebraic topology, but it is a completely elementary parity argument.
@@
The research of P. Soberón was supported by NSF CAREER award no. 2237324, NSF award no. 2054419 and a PSC-CUNY Trad B award.
@@ 1. Introduction
Understanding how one can split families of measures on real spaces into pieces of the same size is a central topic in topological combinatorics [RoldanPensado2022, Zivaljevic2017]. The classic example, known as the ham sandwich theorem, states that given $d$ smooth probability measures on $\mathds{R}^{d}$, there exists a hyperplane simultaneously splitting each measure in half.
@@
A myriad of related results have appeared since then, many of them recover the foodie references (cakes [cccg98-akiyama-perfect], spicy chickens [Karasev:2014gi], pizzas [Barba:2019to], etc.), most of these results are proven with topological methods, which range from fixed point and Borsuk–Ulam type theorems to more advanced techniques in algebraic topology (cohomological index theory, equivariant obstruction theory, spectral sequences, etc). Even splitting measures with families of few hyperplanes can lead to very rich topological statements (see, for instance, the history of the Hadwiger–Grünbaum–Ramos theorem [Blagojevic:2018jc]).
@@ 1.1. Our result
A central instance of this definition is when a finite measure $\mu$ is absolutely continuous with respect to the Lebesgue measure (or more generally, $\mu(h)=0$ for every affine hyperplane $h$), in this case the bisecting condition is simply:
@@
  • The case $d=1$, $n=1$ and any $k$ is the “necklace splitting theorem”. This was originally proved by Hobby and Rice [Hobby:1965bh] and then rediscovered in the 80’s with two new proofs [Goldberg:1985jr, Alon:1985cy]. As $S(k,k)=1$, we don’t impose any new requirement on $k$.
@@ 1.4. Summary
The proof of Theorem 1.1 has three steps. First, we construct a particular family of measures supported in sets of an odd number of points for which the number of bisecting arrangements is exactly $N$. Second, we show that for any deformation of these point sets the parity of the number of bisecting hyperplane arrangements is invariant.
Finally, we show that the result for measures supported on an odd set of points implies the general result.
@@ 2. The ham sandwich theorem revisited
In this section we show how our method give a new and elementary proof of the ham sandwich theorem. This is a particular case of the general proof described in the next sections, but may help the reader build intuition about this approach. Due to the equivalence of the Borsuk–Ulam theorem and the ham sandwich theorem [Karthik2017], this can be used as one of the ingredients of a new convoluted proof of the Borsuk–Ulam theorem.
@@ 3.1. Points in generic position
The proof is analogous to the one for Lemma 3.1. Let $X=X_{1}\cup\dots\cup X_{k}$. Pick a point $p_{i}\in X_{i}$ for each $i=1,\dots,k$, and consider the $l-2$ vectors of the form $p-p_{i}$ for $p\in X_{i}\setminus\{p_{i}\}$. These, together with $B_{L}$, form $d-2$ linearly independent vectors, which gives us the basis for $K$. The $k$ translates of $K$ containing each of $p_{1},\dots,p_{k}$ are the affine spaces we were looking for. ∎
@@ 4. Necessity of conditions
In this case, we construct $(M+1)/2$ pairs of parallel segments as before, and then remove one segment of the final pair. Assume that we have a valid arrangement of hyperplanes that bisects this last segment. If at least two hyperplanes cut this last segment, then we do not have enough hyperplanes to bisect the remaining $(M-1)/2$ pairs of tiny parallel segments. If exactly one hyperplane cuts the last segment, we may assume without loss of generality that is is not one of the hyperplanes orthogonal to $L_{1}$ (otherwise, replace the roles of $L_{1}$ and $L_{2}$ in the rest of the argument). Then, since we still have to use an odd number of hyperplanes orthogonal to $L_{1}$, the argument follows as in case $1$. ∎